# FALCON: Learning Force-Adaptive Humanoid Loco-Manipulation

### v1

@@ A.1 Reward Terms
We adopt the similar reward terms from [17, 11], but introduce some important penalties to ensure the locomotion stability under significant external forces, and other tracking rewards for squat and waist twist. The additioanl reward terms are summerized in Table 5:
@@
Hip pos $\lVert\boldsymbol{q}^{hip}_{roll,pitch}\rVert$2.5
Negative knee joint $\sum_{j}\mathds{1}[q_{j}<q_{j}^{\min}]$1.0
@@
Stance root $|(\boldsymbol{p}_{\text{root}}-\text{mid}(\boldsymbol{p}_{\text{feet}}))^{y}|$5.0
Stand still $\mathds{1}[\text{no contact}]$0.15
Ankle roll $\sum_{j}|q_{j}^{\text{roll}}|$2.0
@@ A.3 Lower-RL-Upper-IK with Force Estimator
We jointly train a 3D force estimator, following a similar approach to [38], using the robot’s proprioception as input $\boldsymbol{s}_{t}^{\mathrm{p}}\triangleq[{\boldsymbol{{q}}_{t-4:t}},{\boldsymbol{\dot{q}}_{t-4:t}},{\boldsymbol{\omega}^{\text{root}}_{t-4:t}},{\boldsymbol{g}_{t-4:t}},\boldsymbol{a}^{l}_{t-5:t-1}].$ As illustrated in Figure 7 (a), the estimator predicts the end-effector forces $\tilde{\boldsymbol{F}}_{t}^{ee}$, which are then concatenated with full-body proprioception and fed into the lower-body RL policy. Meanwhile, the upper-body joint torques with force compensation are computed as $\boldsymbol{\tau}=K_{p}(\boldsymbol{q}^{\text{upper}}_{t}-\boldsymbol{q}^{\text{upper*}}_{t})+K_{d}\boldsymbol{\dot{q}}^{\text{upper}}_{t}+\boldsymbol{J}_{EE}^{T}\tilde{\boldsymbol{F}}_{t}^{ee}.$
@@ A.5 Autonomy Pipeline
We develop an hierarchically autonomous pipeline for tote logistics, leveraging a Motion Capture (MoCap) system to localize positions of the robot and desks. The robot is controlled by a state-machine framework with four states: (1) walking without the tote, (2) picking up the tote in stance, (3) walking with the tote, and (4) dropping off the tote in stance, as illustrated in Fig.9. To estimate the tote’s pose relative to the camera, we use FoundationPose [41], a state-of-the-art method for accurate and reliable 6-DoF pose estimation.
@@ A.5.1 Perception - Pose Estimation
To set up the FoundationPose pipeline [41], we first acquire a high-fidelity 3D model of the industrial tote by performing a raw 3D scan, followed by manual post-processing in a 3D modeling tool. The resulting texture and .obj files serve as inputs to FoundationPose, enabling 6-DoF pose estimation of the tote from the G1 robot’s image stream. Additionally, we predefine grasp points longitudinally on the tote’s surfaces (Fig. 10), which are transformed into the robot base frame using the calibrated extrinsics between the camera and the robot.
@@ A.6 Hardware Limits in Real-World Deployment
During sim-to-real deployment of the policy trained with FALCON, we observe that the humanoid robot struggles to sustain high joint torques over extended periods, often leading to rapid motor overheating—particularly at the wrists, as shown in Figure 11 (a). This significantly limits our ability to perform payload transport exceeding 2kg per arm at its default joint position. In contrast, as shown in Figure 11 (b) and (c), the same policy evaluated in MuJoCo [47], with torque clipping to respect joint limits but without modeling thermal constraints, successfully transports payloads over 3kg per end effector while accurately tracking the linear velocity commanded $-1$ m / s along the x-axis. This highlights a key gap between simulated and real-world actuator endurance.

### v2

@@ A.1 Reward Terms
We adopt the similar reward terms from [he2024omnih2o, cheng2024expressive], but introduce some important penalties to ensure the locomotion stability under significant external forces, and other tracking rewards for squat and waist twist. The additioanl reward terms are summerized in Table 5:
@@
Hip pos $\lVert\boldsymbol{q}^{hip}_{roll,pitch}\rVert$ -2.5
Negative knee joint $\sum_{j}\mathds{1}[q_{j}<q_{j}^{\min}]$ -1.0
@@
Stance root $|(\boldsymbol{p}_{\text{root}}-\text{mid}(\boldsymbol{p}_{\text{feet}}))^{y}|$ -5.0
Stand still $\mathds{1}[\text{no contact}]$ -0.15
Ankle roll $\sum_{j}|q_{j}^{\text{roll}}|$ -2.0
@@ A.3 Lower-RL-Upper-IK with Force Estimator
We jointly train a 3D force estimator, following a similar approach to [portela2024learning], using the robot’s proprioception as input $\boldsymbol{s}_{t}^{\mathrm{p}}\triangleq[{\boldsymbol{{q}}_{t-4:t}},{\boldsymbol{\dot{q}}_{t-4:t}},{\boldsymbol{\omega}^{\text{root}}_{t-4:t}},{\boldsymbol{g}_{t-4:t}},\boldsymbol{a}^{l}_{t-5:t-1}].$ As illustrated in Figure 7 (a), the estimator predicts the end-effector forces $\tilde{\boldsymbol{F}}_{t}^{ee}$, which are then concatenated with full-body proprioception and fed into the lower-body RL policy. Meanwhile, the upper-body joint torques with force compensation are computed as $\boldsymbol{\tau}=K_{p}(\boldsymbol{q}^{\text{upper}}_{t}-\boldsymbol{q}^{\text{upper*}}_{t})+K_{d}\boldsymbol{\dot{q}}^{\text{upper}}_{t}+\boldsymbol{J}_{EE}^{T}\tilde{\boldsymbol{F}}_{t}^{ee}.$
@@ A.5 Autonomy Pipeline
We develop an hierarchically autonomous pipeline for tote logistics, leveraging a Motion Capture (MoCap) system to localize positions of the robot and desks. The robot is controlled by a state-machine framework with four states: (1) walking without the tote, (2) picking up the tote in stance, (3) walking with the tote, and (4) dropping off the tote in stance, as illustrated in Fig.9. To estimate the tote’s pose relative to the camera, we use FoundationPose [wen2024foundationpose], a state-of-the-art method for accurate and reliable 6-DoF pose estimation.
@@ A.5.1 Perception - Pose Estimation
To set up the FoundationPose pipeline [wen2024foundationpose], we first acquire a high-fidelity 3D model of the industrial tote by performing a raw 3D scan, followed by manual post-processing in a 3D modeling tool. The resulting texture and .obj files serve as inputs to FoundationPose, enabling 6-DoF pose estimation of the tote from the G1 robot’s image stream. Additionally, we predefine grasp points longitudinally on the tote’s surfaces (Fig. 10), which are transformed into the robot base frame using the calibrated extrinsics between the camera and the robot.
@@ A.6 Hardware Limits in Real-World Deployment
During sim-to-real deployment of the policy trained with FALCON, we observe that the humanoid robot struggles to sustain high joint torques over extended periods, often leading to rapid motor overheating—particularly at the wrists, as shown in Figure 11 (a). This significantly limits our ability to perform payload transport exceeding 2kg per arm at its default joint position. In contrast, as shown in Figure 11 (b) and (c), the same policy evaluated in MuJoCo [mujoco], with torque clipping to respect joint limits but without modeling thermal constraints, successfully transports payloads over 3kg per end effector while accurately tracking the linear velocity commanded 1-1 m / s along the x-axis. This highlights a key gap between simulated and real-world actuator endurance.